THE TEMPERATURE DEPENDENCE OF THE RAMAN BANDWIDTHS FOR THE HARD MODES OF THE AMMONIUM HALIDES CLOSE TO PHASE TRANSITIONS

\underline{H. YURTSEVEN}, Department of Physics, Middle East Technical University, Ankara-Turkey.